### v1

@@ 3.2 Scene, Object and Trajectory Generation
$F_{\theta}(\mathbf{z}_{k},\mathbf{c})$ is the score function $\nabla_{\mathbf{z}}\log p(\mathbf{z}|\mathbf{c})$ of the conditional distribution at $\mathbf{z}_{k}$ and $F_{\theta}(\mathbf{z}_{k},\mathbf{c}=\mathbf{0})$ is the the score function of the unconditional distribution $p_{\theta}(\mathbf{z})$. $w$ is the CFG guidance scale parameter, $\lambda_{k}$ is an annealed noise schedule parameter, and $\boldsymbol{\epsilon}_{k}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Finally, a 3D scene sample $\mathbf{s}$ is recovered by decode the reverse-diffused sampling latent code $\mathbf{s}=E_{\theta}(\mathbf{z}_{0})$. Fig. 4(c) depicts the sampling procedure.
@@
Training: We train our diffusion-based scene generation model using a dataset that pairs scene geometry with map conditioning. Direct access to dense scene geometry is not available in practice. Instead, we use the state-of-the-art dense geometry reconstruction approach, neural kernel surface reconstruction (NKSR) [28], to recover a pseudo-GT from an input LiDAR sequence. Ground truth annotations are used to remove moving dynamic objects, ensuring our reconstruction contains only the static scene and objects. We then train the auto-encoders for both scene geometry and map layout using reconstruction loss and KL divergence loss: $\min_{\theta}\mathcal{L}_{\mathrm{recon}}+\mathcal{L}_{\mathrm{KL}}$ over real-world examples (Fig. 4 (a)). Our latent diffusion model is trained using the score matching loss function: $\mathcal{L}_{\mathrm{LDM}}=\mathbb{E}_{(\mathbf{z},\mathbf{c}),\boldsymbol{\epsilon},k}\left[\|\boldsymbol{\epsilon}-F_{\theta}(\mathbf{z}_{k},k,\mathbf{c})\|_{2}^{2}\right]$, where $\mathbf{z}_{k}$ is the forward diffused noisy sample at step $k$ (Fig. 4 (b)).
Figure 5: Stochastic raydrop networks for sensory noise simulation, further enhancing realism. We highlighted the raydropped points in red on Masked Range and Masked LiDAR Images above.
Object Generation.
Inspired by the rapid advancements in 3D shape generation, we employ two high-fidelity, 3D object generation frameworks, GET3D [19] and AvatarClip [26], to create dynamic traffic participants.
@@
For each actor $\mathbf{o}_{i}$ in a given layout $\mathcal{I}$, we sample a random variable $\mathbf{z}\sim\mathcal{N}$ and generate the corresponding actor mesh following $\mathbf{o}_{i}=G(\mathbf{z})$, where $G(\cdot)$ is the Generator/Decoder of the chosen generative method. For cars, trucks, and other four-wheeled vehicles, we use GET3D [19], which has demonstrated state-of-the-art and diverse generative results for 3D shapes, including cars. The layout box in $\mathcal{I}$ is used to rescale the shape of each actor, ensuring that the sizes of the generated shape and the input layout are consistent. AvatarClip [26] is used to generate pedestrians conditioned on a SMPL [41] pose and shape parameter $\bf{p}=(\boldsymbol{\theta},\boldsymbol{\beta})$. Furthermore, each generated rigged model is animated with a walking animation from Mixamo [3], ensuring a realistic 3D human walking motion sequence over time.
@@
Together, the generated static world $\mathbf{s}$ and each actor $\mathbf{o}_{i}$ define our 3D world scenario, denoted as $\mathcal{W}$. Fig. 3 depicts the composed scene.
Trajectory Generation.
To simulate a dynamic traffic scenario, we propose a retrieval-augmented generation coupled with a rejective sampling scheme to generate realistic and physically plausible dynamics for each actor and the ego vehicle, turning our 3D scene representation $\mathcal{W}$ to 4D, denoting as $\mathcal{W}_{t}$ for a given time $t$.
For the ego vehicle and each actor, jointly denoted as $\mathcal{P}=\{\tau_{t}\}$, we sampled trajectories from a trajectory bank obtained from Waymo Open dataset [63] and augment them to the scene. This guarantees realistic dynamics as the dataset is obtained from a diverse set of real-world scenarios. To ensure the physical feasibility of the sampled trajectories with respect to our generated scene, we reject those that violate rules of physics, such as collision with the static world, collision between actors, or hovering over the non-mesh area, while making sure to position them correctly above the ground. Around 12.3% of sampled trajectories are retained. This rate is acceptable because resampling is trivial.
@@
Additionally, for more realistic trajectory generation targeting for simulation use-cases, we extend Waymax [20], a data-driven 2D BEV traffic simulator, to control the behaviors of traffic actors in more systematic manners. Given a scenario from the WOMD Dataset [16], we use Waymax to replay ego-vehicle’s and agent’s real-world trajectories, with an additional reactive intelligent driving module [69] that updates each agent’s acceleration to avoid collisions. Since the trajectories are specific to the given real-world scenario, they are guaranteed to be physically plausible, but less diverse than our approach above.
@@
Effectively, this approach renders our world generation to be completely asset-free, end-to-end generative, and thereby temporally consistent, allowing for a realistic, generative, and physics-based simulation without the need for artist-curated [14] or pre-collected assets [46, 74] as in previous lidar simulation method.
3.3 Physics-Informed LiDAR Generation
@@
Given the complete 4D world $\mathcal{W}$ and the poses $\mathcal{P}$, our next step is to generate a realistic LiDAR point cloud corresponding to these conditions. At a high level, we use the poses to compose the scene and objects at each timestep, then perform physics-informed ray casting to obtain purely physically simulated LiDAR as an intermediate result. As a final step, we leverage data-driven conditional sampling to generate the final LiDAR point cloud to simulate real-world LiDAR noises from the clean ray casting LiDAR.
Scene Composition.
@@
We use the Dual Marching Cube method [57] to obtain the 3D mesh of the static world from the generated TSDF volume $\mathbf{s}$. Then, with the trajectories of the ego car and all actors at time $t$, $\tau_{t}$, we transform the 3D mesh to the world coordinates and compose it with $\mathbf{s}$, producing the full world geometry at each time $t$: $\mathcal{W}_{t}=\pi(\mathcal{W},\boldsymbol{\tau}_{t})$ For each vehicle actor, $\pi$ applies a rigid body transformation. For pedestrians, besides the rigid body movement, $\pi$ additionally articulates the human body shape to simulate the animated nonrigid human movement with forward kinematics [7].
Physics-based Ray Casting.
From the ego vehicle position $\tau_{t}$, we perform raycasting by utilizing Open3D [81] to compute ray-triangle intersections against the composed scene $\mathcal{W}_{t}$, obtaining the raycast scan $\overline{\mathbf{x}}_{t}$. To enhance realism of the raycasting to real-world LiDAR, we closely match the LiDAR sensor configuration (elevation angles, azimuth angles, field of view, etc.) with the real-world Lidar, depending on the use-case. For single-frame generation (Sec. 4.3), we follow the Velodyne HDL-64E manual [1] to match with KITTI-360 dataset [39]. For conditional multi-frame generation (Sec. 4.4), we use the calibration information provided with Waymo Open Dataset [63].
Stochastic Raydrop.
@@
Raycasted LiDAR from the generated world appears over-clean without real-world noises due to environmental and sensor noise factors. To address this, inspired by LiDARSim, we have an additional stage that stochastically simulates “raydrop”, where rays do not return to the sensor. For each raycast scan at time $t$, $\overline{\mathbf{x}}_{t}$, we project it onto a 2D spherical range image using polar coordinates. This image uses azimuth and zenith angles to represent coordinates and encodes depth values for each pixel. We predict raydrop probability per pixel on this image using a U-Net architecture [48] supervised by real-world LiDAR scan raydrop masks. Our approach, unlike LiDARSim, requires only a range map, eliminating the need for multiple additional metadata input channels that are only available in real-world data. We also apply a Gumbel sigmoid for random sampling. The application of our stochastic raydrop method produces the final LiDAR scan $\mathbf{x}_{t}$ for each frame, concluding in our complete end-to-end LiDAR video generation process, as shown in Fig. 5.
4 Experiment
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Real ProjectedGAN LiDARGen UltraLiDAR Ours
@@ 4 Experiment
Figure 6: Real KITTI-360 samples vs unconditional samples from the competing methods. UltraLiDAR sample visualizations are directly acquired from their paper. Compared to previous approaches, LidarDM generates samples that feature a greater quantity of more detailed salient objects (e.g., cars, pedestrians), sharper 3D structures (e.g., straight walls), and more realistic road layouts.
4.1 Setup
Datasets.
We evaluate our proposed LidarDM on KITTI-360 [39] and Waymo Open [63] datasets. KITTI-360 is the de facto dataset for evaluating unconditional LiDAR generation methods. The dataset contains nine driving sequences (76,715 samples), where the first sequence is used as a val sequence (11,518 samples) and the last eight are used for training (65,197 samples). The sequences were collected across Karlsruhe, Germany with a 64 Beam LiDAR. However, KITTI-360 does not provide detailed BEV HD map information limiting its applications in conditional models. Waymo Open [63] is a dataset containing 1048 sequences with 158,000 training and 29,700 validation frames. A detailed HD map allows us to train conditional diffusion models. The HD map is in a vector format storing edges of map objects. We preprocess the map by centering it on each LiDAR Frame and rasterizing it into a segmentation map. The dimensions of the map tensor are L×$\times$×W×$\times$×M (length×$\times$×width×$\times$×map classes). The map has 5 classes (lane markings, road lines, edges, crosswalks, driveways).
Training Details.
We train our models using four Nvidia A100 40GB GPUs. We use the AdaM optimizer with a learning rate of 1e-4 for the VAE and 1e-5 for the diffusion U-Net, with a cosine decay schedule.
Model Details.
The diffusion UNet, which is conditioned on layout and SDF latents, has 5 ResNet blocks with 2$\times$ downsampling with channels of
{128,128,256,512,512}. The SDF VAE has 4 ResNet blocks with channels of {448,640,896,1280}. The Map VAE has ResNet down/upsample blocks with channels of {64,64,128,256,512}, and it is trained independently with X-Entropy and KL regularization.
4.2 Baselines
Unconditional Generation.
LiDARVAE, LiDARGAN, ProjectedGAN, and LiDARGen are baselines that perform generation in the range image representation. UltraLiDAR performs generation in the BEV voxel space. To provide a fair comparison, we follow UltraLiDAR and evaluate MMD and JSD on a histogram of voxel occupancy instead of voxel density [72]. In addition, UltraLiDAR does not provide samples so we use their reported numbers in their original paper.
Temporal Coherency.
We are the first to attempt the task of sequential LiDAR generation and thus no previous models exist for comparison. Nonetheless, we implement a sequence diffusion baseline inspired by recent work in video generation. We believe this approach is the most straightforward initial approach.
Concretely, we train a VAE to encode individual LiDAR frames. This has been shown previously [72] to be powerful and effective. Next, we train a diffusion model to directly denoise multiple (i.e., 5) LiDAR frames at once. Visually, this approach yields decent temporal consistency.
4.3 Unconditional Single-Frame Generation
We first validate our model architecture design and showcase our model’s generative capability by directly comparing against previous LiDAR generation models in unconditional generation on KITTI-360 [39].
Based on the results in Table 1, BEV models (such as ours or UltraLiDAR) perform best in top-down layout quality (as reflected by MMD and JSD) compared to range image models. The close performance gap compared to UltraLiDAR can be explained by the fact that UltraLiDAR was directly trained on the task of modeling single LiDAR scans which the benchmark evaluates. In addition, the BEV voxel grid representation offers large flexibility in generating physically implausible LiDAR readings that have the potential accurately match ground truth data in histogram metrics. We also show qualitative comparisons against the baselines in Fig. 6.
@@ 4.3 Unconditional Single-Frame Generation
LiDAR VAE [8] $8.53\times 10^{-4}$ $0.267$
LiDAR GAN [8] $8.95\times 10^{-4}$ $0.243$
ProjectedGAN [56] $7.07\times 10^{-4}$ $0.201$
LidarGen [83] $2.95\times 10^{-4}$ $0.136$
UltraLidar [72] $\mathbf{9.67\times 10^{-5}}$ $0.132$
LidarDM (Ours) ${1.67\times 10^{-4}}$ $\mathbf{0.119}$
Table 1: Qualitative results for unconditional generation on KITTI-360 dataset. (  best,   second best,   third best)
4.4 Map-conditioned Multi-Frame Generation
@@
4.4.1 Consistent Video Generation
Method Total ICP Energy [m] ($\downarrow$) Average ICP Energy ($\downarrow$) Outlier Percentage ($\downarrow$) Chamfer Distance [m] ($\downarrow$)
Sequence Diffusion 3616.58 0.078 20.56% 0.39
LidarDM (Ours) 916.94 0.014 7.12% 0.17
@@ 4.4.1 Consistent Video Generation
Table 2: Temporal consistency. Outlier percentage uses distance threshold $\tau=0.5m$.
@@
One of our key contributions is the temporal consistency of the sequential LiDAR generation. To evaluate this, we first use ICP alignment to calculate a relative transformation between consecutive generated frames. We then exploit LiDAR’s 3D nature and define an average point-to-plane energy over a sequence of LiDAR scans as our quantitative metrics, following this equation: $E=\dfrac{1}{T}\sum_{t=1}^{T}\texttt{point2plane}(x_{t},x_{t-1})$ where point2plane represents the point-to-plane distance [42], and $x_{t}$ indicates the LiDAR scan at time $t$. Intuitively, $E$ is prone to higher values from dynamic objects, but it is still a valuable metrics to determine if the general geometry of the 3D scene is preserved over time as the majority of the environment is static. To further evaluate the geometric consistency of the generative LiDAR sequence, we also measure the outlier point ratio, defined as the percentage of points with the point2plane distance larger than a certain threshold $\tau$. Table 2 shows our quantitative results, where we beat the baseline in both metrics by a notable margin. These results clearly show that our LidarDM is capable of generating temporally-consistent LiDAR sequences.
[…]
Real 59.7 81.1%
LidarDM 56.4
Table 3: Real2Sim: Detector [75] trained on real data can be evaluated on LidarDM-generated data, showing strong agreement with its real counterpart, suggesting its potential for simulation.
[…]
35k Real $58.2$
35k Real + 70k LidarDM $61.3$
Table 4: Data augmentation: LidarDM-generated data can enhance training for detectors on real-world data, suggesting its potential to improve the perception module without expensive data annotation.
Figure 7: Qualitative results of map-conditioned sequence generation on 2 Waymax [20] map sequences. We also showcase the corresponding accumulated point cloud to hightlight the temporal consistency of LidarDM.
4.4.2 Layout-aware LiDAR Generation

### v2

@@ III-B1 Scene Generation
$F_{\theta}(\mathbf{z}_{k},\mathbf{c})$ is the score function $\nabla_{\mathbf{z}}\log p(\mathbf{z}|\mathbf{c})$ of the conditional distribution at $\mathbf{z}_{k}$ and $F_{\theta}(\mathbf{z}_{k})=F_{\theta}(\mathbf{z}_{k},\mathbf{c}=\mathbf{0})$ is the the score function of the unconditional distribution $p_{\theta}(\mathbf{z})$. $w$ is the CFG guidance scale parameter, $\lambda_{k}$ is an annealed noise schedule parameter, and $\boldsymbol{\epsilon}_{k}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Finally, a 3D scene sample $\mathbf{s}$ is recovered by decoding the reverse-diffused sampling latent code $\mathbf{s}=E_{\theta}(\mathbf{z}_{0})$.
@@
Training: We train our diffusion-based scene generation model using a dataset that pairs scene geometry with map conditioning. Direct access to dense scene geometry is not available in practice. Instead, we use NKSR [48], a textured 3D reconstruction method, to recover a pseudo-GT from an input LiDAR sequence. The recovered mesh’s texture encodes intensity. Ground truth annotations are used to remove moving dynamic objects, ensuring our reconstruction contains only the static scene and objects. We then train the auto-encoders for both scene geometry and map layout using reconstruction loss and KL divergence loss: $\min_{\theta}\mathcal{L}_{\mathrm{recon}}+\mathcal{L}_{\mathrm{KL}}$ over real-world examples. Our latent diffusion model is trained using the score matching loss function: $\mathcal{L}_{\mathrm{LDM}}=\mathbb{E}_{(\mathbf{z},\mathbf{c}),\boldsymbol{\epsilon},k}\left[\|\boldsymbol{\epsilon}-F_{\theta}(\mathbf{z}_{k},k,\mathbf{c})\|_{2}^{2}\right]$, where $\mathbf{z}_{k}$ is the forward diffused noisy sample at step $k$ .
III-B2 Object Generation
@@
We employ two object generation frameworks, GET3D [5] and AvatarClip [49], to create dynamic traffic participants. For each actor $\mathbf{o}_{i}$ in a given layout $\mathcal{I}$, we sample a random variable $\mathbf{z}\sim\mathcal{N}$ and generate the corresponding actor mesh following $\mathbf{o}_{i}=G(\mathbf{z})$, where $G(\cdot)$ is the generator/decoder of the chosen generative method. We use GET3D to generate vehicles and then rescale them to properly fit within a target bounding box of the layout. For pedestrian generation, we utilize AvatarClip [49], which is conditioned on a SMPL [50] pose and shape parameter $\bf{p}=(\boldsymbol{\theta},\boldsymbol{\beta})$. We use Mixamo [51] to animate the rigged model, ensuring realistic 4D human walking motion.
@@
Together, the generated static world $\mathbf{s}$ and each actor $\mathbf{o}_{i}$ define our 3D world, denoted as $\mathcal{W}$, as depicted in Fig. 3
III-B3 Trajectory Generation
@@
We extend Waymax [45], a data-driven 2D BEV traffic simulator, to control the behaviors of traffic actors in more systematic manners. Given a scenario from the WOMD Dataset [52], we use Waymax to replay ego-vehicle’s and agent’s real-world trajectories, with an additional reactive Intelligent Driver Model [53] that updates each agent’s acceleration to avoid collisions. For unconditional generation, we sample trajectories from a trajectory bank obtained from Waymo Open dataset [54]. We employ heuristics to ensure physical feasibility (no hovering) and that no collisions occur (through agent-agent or agent-scene collisions). Around 12.3% of sampled trajectories are retained, which is acceptable because resampling is trivial.
@@
This approach renders our world generation to be completely asset-free, end-to-end generative, and temporally consistent, allowing for a realistic and physics-based simulation without the need for artist-curated [32] or pre-collected assets [13, 35] as in previous LiDAR simulation methods.
III-C Physics-Informed LiDAR Generation
@@
Given the complete 4D world $\mathcal{W}$ and the poses $\mathcal{P}$, our next step is to generate a realistic LiDAR point cloud corresponding to these conditions. At a high level, we use the poses to compose the scene and objects at each timestep, then perform physics-informed ray casting to obtain purely physically simulated LiDAR as an intermediate result. We leverage data-driven conditional sampling to generate the final point cloud to simulate real-world LiDAR noises.
III-C1 Scene Composition
@@
We use Dual Marching Cube  [55] to obtain the 3D mesh of the static world from the TSDF channel of the generated $\mathbf{s}$. Then, using our generated actor trajectories, we transform all 3D agents’ meshes to the world coordinates and compose it with $\mathbf{s}$ using ego poses and all actors actor poses at time $t$, $\tau_{t}$, producing the full world geometry at each time $t$: $\mathcal{W}_{t}=\pi(\mathcal{W},\boldsymbol{\tau}_{t})$. For vehicles, $\pi$ applies a rigid transform. For pedestrians, $\pi$ applies a rigid transform and articulates the human body shape to simulate animated movement with forward kinematics [56].
III-C2 Physics-based Ray Casting
LiDAR sensors acquire a 3D point cloud by shooting beams of light from the sensor into the scene. Raycasting simulates this process for a single beam by calculating the $\{x,y,z\}$ coordinate point where a beam would intersect with the scene based on a given ego vehicle position $\tau_{t}$, the composed scene $\mathcal{W}_{t}$, and the beam’s elevation ($\theta$) and azimuth ($\phi$) angles. LiDARs typically have multiple beams that spin to generate a single 3D point cloud. Therefore, the process is repeated for each beam based on the virtual sensor configuration which we match with the real-world LiDARs based on provided KITTI HDL-64E [57] and Waymo specifications.
III-C3 Intensity Modeling
For each point $\overline{\mathbf{x}}^{s}_{it}\in\overline{\mathbf{x}}_{t}$ that lies on the background scene, we query the intensity channel of the volume $\mathbf{s}$ for the intensity of the closest vertex. For intensity values of dynamic objects, we opt for a physical-based approach following Lambert’s Cosine Law. Namely, for each point $\overline{\mathbf{x}}^{o}_{it}\in\overline{\mathbf{x}}_{t}$ that lies on the dynamic object, $intensity(\overline{\mathbf{x}}^{o}_{it})=\alpha+\beta\left(I_{it}\cdot N^{o}_{t}\right)$ where $I_{it}$ denotes the vector from the lidar sensor to $\overline{\mathbf{x}}^{o}_{it}$, $N^{o}_{t}$ denotes the normal vector of the object at $\overline{\mathbf{x}}^{o}_{it}$, and $\alpha,\beta$ are constants that have been hand-tuned to match the empirical intensity distribution of dynamic objects in Waymo Dataset.
III-C4 Stochastic Raydrop
@@
Raycasted LiDAR from the generated world appears overly clean, without real-world environmental and sensor noise. Inspired by LiDARSim, we stochastically simulates “raydrop”, where rays do not return to the sensor. For each raycast scan at time $t$, $\overline{\mathbf{x}}_{t}$, we project it onto a 2D spherical range image and predict raydrop probability per pixel on this image using a U-Net architecture [58], supervised by real-world LiDAR scan raydrop masks. Our approach, unlike LiDARSim, requires only a range map, eliminating the need for multiple additional metadata input channels that are only available in real-world data. We sample from this mask with a Gumbel sigmoid to produce the final LiDAR scan $\mathbf{x}_{t}$ for each frame, concluding the end-to-end LiDAR video generation process.
IV Experiment
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Real ProjectedGAN LiDARGen UltraLiDAR R2DM Ours
@@ IV Experiment
Figure 4: Real KITTI-360 samples vs unconditional samples from the competing methods. LidarDM generates samples that feature more detailed salient objects (e.g., cars, pedestrians), sharper 3D structures (e.g., walls), and realistic road layouts.
IV-A Setup
IV-A1 Datasets
@@
IV-A2 Training Details
We train our models for 48 hours using four Nvidia A100 40GB GPUs. We use the Adam optimizer with a learning rate of 1e-4 for the VAE and 1e-5 for the diffusion U-Net, with a cosine decay schedule.
IV-A3 Model Details
The latent diffusion is a UNet with 5 ResNet blocks, featuring channels of 128, 128, 256, 512, 512. The SDF VAE has 4 ResNet blocks with channels of 448, 640, 896, 1280. The Map VAE has ResNet down/upsample blocks with channels of 64, 64, 128, 256, 512.
IV-B Baselines
IV-B1 Unconditional Generation
LiDARVAE, LiDARGAN, ProjectedGAN, LiDARGen, and R2DM are baselines that use range image representation whereas UltraLiDAR uses BEV voxel space. For fair comparison, we follow UltraLiDAR and evaluate MMD and JSD on a histogram of voxel occupancy instead of voxel density [29]. [^1] LiDM has random yaw rotations in their samples that can cause unfair comparison (confirmed with authors). We will include once they provide rotation-free samples
IV-B2 Temporal Coherency
We are the first to attempt the task of sequential LiDAR generation and thus no previous models exist for comparison. Nonetheless, we implement a sequence diffusion baseline inspired by recent work in video generation. Concretely, we train a VAE to encode individual LiDAR frames. This has been shown previously [29] to be effective. Next, we train a diffusion model to directly denoise multiple (i.e., 5) LiDAR frames at once.
IV-C Unconditional Single-Frame Generation
We first validate our model architecture design and showcase our model’s generative capability by directly comparing against previous LiDAR generation models in unconditional generation (without using HD maps) on KITTI-360 [59]. Based on the results in Table I, BEV models (ours or UltraLiDAR) perform best compared to range image models. Note that UltraLiDAR was directly trained on the task of modeling single LiDAR scans which the benchmark evaluates, which explains the performance gap compared to ours. We also show qualitative comparisons against the baselines in Fig. 4.
[…]
LiDAR VAE [30] 8.538.53e4-4 $0.267$
LiDAR GAN [30] 8.958.95e4-4 $0.243$
ProjectedGAN [60] 7.077.07e4-4 $0.201$
LidarGen [28] 2.952.95e4-4 $0.136$
UltraLidar [29] 9.679.67e5-5 $0.132$
R2DM [44] 3.603.60e4-4 $0.148$
LidarDM (Ours) 1.671.67e4-4 $\mathbf{0.119}$
TABLE I: Qualitative results for unconditional generation on KITTI-360 dataset. (Int: Intensity, Con: Controllability, Tem: Temporal Consistency) (  best,   second best,   third best)
IV-D Map-conditioned Multi-Frame Generation
@@
IV-D1 Consistent 4D Generation
Metrics Sequence Diffusion LidarDM
Total ICP Energy [m] ($\downarrow$) 3616.58 916.94
Average ICP Energy ($\downarrow$) 0.078 0.014
Outlier Percentage ($\downarrow$) 20.56% 7.12%
Chamfer Distance [m] ($\downarrow$) 0.39 0.17
@@ IV-D1 Consistent 4D Generation
TABLE II: Temporal consistency. Outlier percentage uses distance threshold $\tau=0.5m$.
@@
One of our key contributions is the temporal consistency of the sequential LiDAR generation. To evaluate this, we first use ICP alignment to calculate a relative transformation between consecutive generated frames. We define an average point-to-plane energy over a sequence of LiDAR scans as our quantitative metrics, following this equation: $E=\dfrac{1}{T}\sum_{t=1}^{T}\texttt{point2plane}(x_{t},x_{t-1})$ where point2plane represents the point-to-plane distance [61], and $x_{t}$ indicates the LiDAR scan at time $t$. Intuitively, $E$ is prone to higher values from dynamic objects, but it is still a valuable metric to determine if the general scene geometry is preserved over time. To further evaluate the geometric consistency , we also measure the outlier point ratio, defined as the percentage of points with the point2plane distance larger than a certain threshold $\tau$. Table II shows our quantitative results, where we beat the baseline in both metrics by a notable margin, clearly demonstrating LidarDM’s temporal consistency.
Figure 5: Qualitative results of map-conditioned generation. Accumulated LiDAR shows temporal and layout consistency.
IV-D2 Layout-aware LiDAR Generation